OpenStudy (anonymous):

Solve for x: 3^(x+2)=27^(2x)

OpenStudy (anonymous):

\[3^{x+2}=3^{3(2x)}\]\[x+2=6x\]\[2=5x\]\[x=2/5\]

OpenStudy (anonymous):

there is a theorem \[b^u = b^v \iff u=v\]that i used to equate the exponents